Adult tickets for a play cost $9 and child tickets cost $8. If there were 23 people at a performance and the theater collected $193 from ticket sales, how many children attended the play?
Answer by jai_kos(139)

Let the Adults be “x”
And Children be “y”
Given 9x + 8y = 193 ------------(1)
x +y = 23 ------------(2)
Now solve the above equations,
Consider the second equation,
x + y = 23
x = 23 – y ------------(3)
Now substitute equation (3) in equation (1), we get
9(23 –y)+ 8y = 193
207 – 9y + 8y = 193
207 – y = 193
207 -193 = y
14 = y
Therefore the number of the children attended the play are 14.